OpenStudy (anonymous):

Given: Pipe 1 d=1 in (diameter) v= 2ft/s (average water flow speed) The equation sets up like the following. Q=Volume flow rate=(average velocity)*(Cross-sectional area of flow) So I get (2ft/s)*(pi/4)(1/12 ft)^2 which would give me a flow rate in terms of (ft/s)^3 Is this set up right? And why/what is pi/4 accounting for?

OpenStudy (anonymous):

\[2\frac{ ft }{s} * \frac{ \pi }{ 4 } * (1 inch)^2= \frac{ ft^3 }{ s }\]

OpenStudy (anonymous):

lol any ideas?